OpenStudy (anonymous):

tom started a business with 500 dollars. 4 months later tina joined him and paid 500 dollars. three months later Raven joined them and paid 500 dollars. at the end of the year the profits were 625 dollars. what does each one get if the profit was given to each depending on how long each one invested

OpenStudy (anonymous):

profit per month: 625/12 = 52.0833$ ..... tom: 4(52.0833)+3(52.0833/2)+5(52.0833/3) = 373.26$ tina: 3(52.0833/2)+5(52.0833/3) = 164.93$ raven: 5(52.0833/3) = 86.8055$